how about this one
2x^2 = 6x

Answer by fractalier(6550)

From 2x^2 = 6x, subtract 6x from both sides, factor and solve, we get
2x^2 - 6x = 0
2x(x - 3) = 0
and thus
x = 0 or x = 3